My homework says "Show that the number is a rationsl by righting it in a/b form. Then give the multiplication inverse and the additive inverse of the number" how would you do that?
1. -4/7
2. 8 2/5
3. -1 1/2

Found 2 solutions by Theo, richwmiller:
Answer by Theo(13342)

the product of a number and it's multiplicative inverse is equal to 1.

see http://www.mathwords.com/m/multiplicative_inverse_of_a_number.htm

the sum of a number and it's additive inverse is equal to 0.

see http://www.mathwords.com/a/additive_inverse_number.htm

to find the multiplicative inverse of a number, you take it's reciprocal.

you take it's reciprocal by dividing the number into 1.

to find the additive inverse of a number, you take it's opposite.

to find it's opposite, you reverse the sign.

example:

multiplicative inverse of 5 is 1/5 because 1/5 * 5 = 1

additive inverse of 5 is -5 because 5 + (-5) = 0

problem number 1:

number is -4/7

multiplicative inverse is -7/4 because -(4/7) * (-7/4) = 1

we got that by taking (-4/7) and dividing it into 1 to get:

1/(-4/7)

if we multiply numerator and denominator of that expression by 7/7, we get:

7/-4 = (-7/4)

additive inverse is 4/7 because -(4/7) + (4/7) = 1

problem number 2:

number is 8 + 2/5

convert this number to an improper fraction to get:

40/5 + 2/5 = 42/5

multiplicative inverse is 5/42 because 5/42 * 42/5 = 1

additive inverse is -42/5 because 42/5 + (-42/5) = 0

problem number 3:

-1 1/2 is the same as -1 - 1/2.

convert to improper fraction to get:

-2/2 - 1/2 = -3/2

multiplicative inverse is -2/3 because -(3/2) * (-(2/3) = 1

additive inverse is 2/3 because -2/3 + 2/3 = 0

note:

improper fraction is a fraction in a/b form where the numerator is greater than the denominator.


Answer by richwmiller(17219)

If this is true then you should throw out your book.
My homework says "Show that the number is a rational by righting it in a/b form.
Righting means to fix something that is wrong.
Writing is to put down on paper.